# Exploring the flow harmonic correlations via multi-particle symmetric and asymmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV

### v2

@@ III.3.2 Six-particle asymmetric cumulants
${\rm AC}_{2,1}(m,n)$ have measured by the ALICE collaboration for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [48]. In Fig.[7], we should the centrality dependence of ${\rm AC}_{2,1}(m,n)$, multiplied by $\langle N_{part}\rangle$ to scale out the trivial dilution of correlation with the increase in the number of multiplets, for Au+Au at $\sqrt{s_{NN}}=200\ GeV$. The magnitudes of $\langle N_{part}\rangle{\rm AC}_{2,1}(2,3)$ and $\langle N_{part}\rangle{\rm AC}_{2,1}(2,4)$ increase with centrality. The trends are similar to the results by the ALICE collaboration. As for symmetric cumulants, negative values of $AC(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as $AC(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. We compute these observables for the three sets of ensembles mentioned in Table  2 and observe that both ${\rm AC}_{2,1}(2,3)$ and ${\rm AC}_{2,1}(2,4)$ are suppressed by shear and bulk viscosities.
@@
Figure 7: Asymmetric cumulant vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm AC}_{2,1}(2,3)$ is consistently negative while as ${\rm AC}_{2,1}(2,4)$ is consistently positive.
@@
In Fig.[8], we have compared the normalized asymmetric cumulant in coordinate and momentum space, as computed using Eq.[8] and the corresponding equation for the initial state eccentricities (Eq.[11]). Again, if only eccentricity drives $v_{n}$, we can expect the ${\rm NAC}(m,n)$ in the final state to be equal to the ${\rm NAC}(m,n)$ in the initial state. Fig. [8] illustrates that the initial negative correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$ primarily causes the observed negative correlation between $v_{2}$ and $v_{3}$. However, unlike symmetric cumulants, here, the initial state correlations are smaller in the peripheral collisions. The correlation between $\varepsilon_{2}$ and $\varepsilon_{4}$ is weaker than the correlation observed between $v_{2}$ and $v_{4}$. The contribution to anisotropic flow $v_{4}$ comes not only from the linear response of the system to $\varepsilon_{4}$, but also from a contribution proportional to $\varepsilon_{2}^{2}$. As collisions become more peripheral, the difference between ${\rm NAC}(2,4)$ in the final state and the initial state increases. This is likely due to $\varepsilon_{2}$ playing a more significant role in $v_{4}$.
Figure 8: Normalized asymmetric cumulants vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm NAC}_{2,1}(2,3)$ is consistently negative while as ${\rm NAC}_{2,1}(2,4)$ is consistently positive.
III.4 Conclusions
@@
We have studied correlators that measure the correlations between flow harmonics of varying orders, specifically four and six-particle correlations, including both symmetric and asymmetric cumulants as functions of centrality at midrapidity in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. These observables are very sensitive to shear and bulk viscosities, resonance decay, and hadronic interactions. A single parameter set of transport coefficients is unable to explain $SC(2,3)$ and $SC(2,4)$ simultaneously.
@@
III.5 Acknowledgements

### v3

@@ IV.3.2 Six-particle asymmetric cumulants
${\rm AC}_{2,1}(m,n)$ have measured by the ALICE collaboration for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [50]. In Fig.[6], we show the centrality dependence of ${\rm AC}_{2,1}(m,n)$, multiplied by $\langle N_{part}\rangle$ to scale out the trivial dilution of correlation with the increase in the number of multiplets, for Au+Au at $\sqrt{s_{NN}}=200\ GeV$. The magnitudes of $\langle N_{part}\rangle{\rm AC}_{2,1}(2,3)$ and $\langle N_{part}\rangle{\rm AC}_{2,1}(2,4)$ increase with centrality. The trends are similar to the results by the ALICE collaboration. As for symmetric cumulants, negative values of $AC(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as $AC(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. We compute these observables for the three sets of ensembles mentioned in Table  2 and observe that both ${\rm AC}_{2,1}(2,3)$ and ${\rm AC}_{2,1}(2,4)$ are suppressed by shear and bulk viscosities. As discussed earlier, the linear component in $v_{4}$ is damped more significantly than the nonlinear term and shows a weaker correlation with $v_{2}$. Thus, as viscosity increases, the correlation between $v_{4}$ and $v_{2}$ is also enhanced.
@@
Figure 6: (Normalized) Asymmetric cumulant vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm AC}_{2,1}(2,3)$ is consistently negative while as ${\rm AC}_{2,1}(2,4)$ is consistently positive.
@@
In Fig.[6], we have compared the normalized asymmetric cumulant in coordinate and momentum space, as computed using Eq.[8] and the corresponding equation for the initial state eccentricities (Eq.[11]). Again, if only eccentricity drives $v_{n}$, we can expect the ${\rm NAC}_{2,1}(m,n)$ in the final state to be equal to the ${\rm NAC}_{2,1}(m,n)$ in the initial state. Fig. [6] illustrates that the initial negative correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$ primarily causes the observed negative correlation between $v_{2}$ and $v_{3}$. However, unlike symmetric cumulants, here, the initial state correlations are smaller in the peripheral collisions. The correlation between $\varepsilon_{2}$ and $\varepsilon_{4}$ is weaker than the correlation observed between $v_{2}$ and $v_{4}$. The contribution to anisotropic flow $v_{4}$ comes not only from the linear response of the system to $\varepsilon_{4}$, but also from a contribution proportional to $\varepsilon_{2}^{2}$. As collisions become more peripheral, the difference between ${\rm NAC}(2,4)$ in the final state and the initial state increases. This is likely due to $\varepsilon_{2}$ playing a more significant role in $v_{4}$.
Currently, STAR has not yet measured asymmetric cumulants. We conducted a comparison between our findings and the Pb+Pb 5.02 TeV results from ALICE (See Appendix Fig.[7] and Fig.[8]). The centrality dependence of the observables we discussed in this paper shows similarities between ALICE and STAR data. The ${\rm SC}(m,n)$ values are about an order of magnitude larger at ALICE than at STAR. Similarly, the ${\rm AC}_{21}(m,n)$ values are 3 to 5 times larger at ALICE compared to STAR. However, the similarities in the magnitude and centrality dependence of the normalized cumulants are noteworthy. This comparison indicates minimal energy dependence for the normalized cumulants.
IV.4 Conclusions
@@
We have studied correlators that measure the correlations between flow harmonics of varying orders, specifically four and six-particle correlations, including both symmetric and asymmetric cumulants as functions of centrality at midrapidity in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. These observables are very sensitive to shear and bulk viscosities, resonance decay, and hadronic interactions. However, this magnified sensitivity can be attributed to the properties of anisotropic flow, particularly as the cumulants incorporate higher powers of flow. We then checked the sensitivities of the normalized cumulants. We found that ${\rm NSC}(2,3)$ and ${\rm NAC}_{2,1}(2,3)$ are insensitive to both the hydro model parameters (shear and bulk viscosity) and late-stage hadronic interactions. This strongly supports their effectiveness as reliable tools for constraining the initial state of the system’s evolution. Conversely, ${\rm NSC}(2,4)$ and ${\rm NAC}_{2,1}(2,4)$ show considerable sensitivity to these stages, allowing them to effectively constrain hydrodynamic and transport models.
A single parameter set of transport coefficients is unable to explain $SC(2,3)$ and $SC(2,4)$ simultaneously.
@@
The values of ${\rm SC}(m,n)$ and ${\rm AC}_{21}(m,n)$ at ALICE are larger—by factors ranging from 3 to 8—compared to those at STAR. However, it is important to note the similarities in the magnitude and centrality dependence of the normalized cumulants. This comparison indicates that there is minimal energy dependence for the normalized cumulants.
IV.5 Acknowledgements